OpenStudy (anonymous):

Solve for the variable 15 + n = 24

OpenStudy (anonymous):

n=24-15

OpenStudy (anonymous):

9

OpenStudy (anonymous):

n = 9 How: 24 - 15 = 9 So 15 + 9 = 24